Trace the curve y=x³-x

1
Expert's answer
2022-02-01T12:18:52-0500

1) We calculate the roots:

"x^3 -x = 0, \;\; x(x^2-1) = 0, \\ 
x = 0 \;\;\text{or} \;\; x = 1 \;\;\text{or} \;\; x = -1 ."

2) Next, we shall determine maxima and minima:

"(x^3-x)' = 3x^2 - 1,\\
3x^2 - 1 = 0,\\
x = \pm \dfrac{1}{\sqrt{3}}."

At "x = -\dfrac{1}{\sqrt{3}}" the function "(x^3-x)'" changes sign from + to -, so it is the point of maximum. Another point is the point of minimum.

At "x = -\dfrac{1}{\sqrt{3}}" the function "x^3-x" obtains value of "\dfrac{2}{3\sqrt{3}}," at "x = +\dfrac{1}{\sqrt{3}}" the function "-\dfrac{2}{3\sqrt{3}}."

3) Next, we'll determine if the function is odd or even.

"f(x) = x^3 - x, \\ f(-x) = (-x)^3 - (-x) = -x^3 + x = -f(x)," so the function is odd.